# Stronger Constraints on Primordial Black Holes as Dark Matter Derived from the Thermal Evolution of the Intergalactic Medium over the Last Twelve Billion Years

### v1

@@ Abstract
Primordial black holes (PBHs) have been explored as potential dark matter candidates, with various astrophysical observations placing upper limits on the fraction $f_{\mathrm{PBH}}$ of dark matter in the form of PBHs. However, a largely underutilized probe of PBH abundance is the temperature of the intergalactic medium (IGM), inferred from the thermal broadening of absorption lines in the Lyman-$\alpha$ forest of quasar spectra. PBHs inject energy into the IGM via Hawking radiation, altering its thermal evolution. In this work, we constrain this energy injection by self-consistently modeling its interplay with the cosmological ultraviolet background from galaxies and supermassive black holes. Leveraging IGM temperature measurements spanning the past twelve billion years ($z\sim 0$ to $6$), we derive one of the most stringent constraints on PBH-induced heating from light PBHs within the mass range $10^{15}$$10^{17}$ g. Specifically, for $M_{\mathrm{PBH}}=10^{16}$ g, we find $f_{\mathrm{PBH}}<5\times 10^{-5}$ at 95% confidence, with the bound scaling approximately as $M_{\mathrm{PBH}}^{4}$ at other masses. Our inclusion of helium reionization and low-redshift temperature measurements strengthens previous IGM-based PBH constraints by an order of magnitude or more. Compared to other existing limits, our result is among the strongest, second only to the constraints from the 511 keV line from the Galactic Centre, but with distinct systematics. More broadly, this study highlights the IGM thermal history as a powerful and independent probe of beyond-standard-model physics.
@@ I Introduction
It is now known that the IGM at redshifts $z<6$ is in thermal equilibrium with a cosmological radiation background created by star-forming galaxies and accreting supermassive black holes. Under certain conditions, this radiative background imparts the IGM a power-law temperature-density relation, usually written as [1],
@@
Fortunately, the temperature of the IGM can also be measured out at least to $z\sim 6$. Most of these measurements use observations of the Lyman-$\alpha$ forest to infer the IGM temperature. The Lyman-$\alpha$ forest is a series of absorption lines observed in the spectra of distant luminous quasars and serves as an excellent tracer of the intervening diffuse gas (see Refs. [2, 3] for reviews). The individual absorption lines in the Lyman-$\alpha$ forest are caused by the neutral hydrogen residing in the IGM. Consequently, the widths of these lines are sensitive to the thermal state of the IGM. The thermal motion of neutral hydrogen introduces a broadening in the width of the lines, expressed as $\Delta\nu=\nu_{\mathrm{Ly}\alpha}b_{\mathrm{th}}/c$, where $c$ is the speed of light, $\nu_{\mathrm{Ly}\alpha}$ is the Lyman-$\alpha$ absorption frequency, and $b_{\mathrm{th}}=\sqrt{2k_{\mathrm{B}}T/m_{\mathrm{H}}}$ is the Doppler parameter associated with the thermal motion with $k_{\mathrm{B}}$ being the Boltzmann’s constant, $m_{\mathrm{H}}$ the mass of the hydrogen atom, and $T$ the temperature of the IGM. By directly measuring the Doppler parameter of individual absorption lines in the Lyman-$\alpha$ forest, or by indirectly measuring it from other statistics sensitive to thermal broadening, it is therefore possible to infer the temperature of the IGM [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. This approach has revealed that $T_{\mathrm{IGM}}$ is of the order of $\sim 10^{4}$ K in the twelve billion years since $z\sim 6$, with some evolution. At $z\sim 6$, the temperature is around $11,500$ K at $z\sim 5.8$ and gradually declines to $\sim 7,300$ K by $z\sim 4.6$. Due to helium reionization at $z\sim 3$, the temperature peaks at approximately $14,750$ K before cooling again to $\sim 4,700$ K at $z\sim 0.06$. The typical uncertainties in these measurements range from $10\%$ to $20\%$.
@@
Exotic energy injection processes can significantly affect the thermal and ionization history of the IGM. Examples include dark matter decay or annihilation, dark matter-baryon interactions, ultralight dark photon dark matter, and light primordial black holes (PBHs). Nonetheless, measurements of the IGM temperature from the Lyman-$\alpha$ forest as a probe of these processes have largely been overlooked so far. Indeed, while the Lyman-$\alpha$ forest has been used to constrain exotic processes, these constraints have primarily relied on measurements of the Lyman-$\alpha$ transmission power spectrum, which probes small-scale cosmological matter density fluctuations; these constraints have largely disregarded the information contained in the Lyman-$\alpha$ forest about the thermal state of the IGM. Examples include warm dark matter [23, 24, 25, 26, 27, 28, 29, 30], sterile neutrinos [23, 24, 27, 31], fuzzy dark matter [32, 33], ultralight axions [34], primodial black holes (PBHs) with masses around $100$ M [35], and isocurvature fluctuations [36]. The potential of IGM temperature as a diagnostic tool for exotic physics remains largely untapped [37, 38, 39, 40].
@@
In this study, we revisit the use of IGM temperature measurements from the Lyman-$\alpha$ forest to place constraints on light PBHs. Such light PBHs in the mass range of $10^{15}$ to $10^{17}\,\mathrm{g}$ have recently gained attention as potential dark matter candidates. Through Hawking evaporation, they emit particles that can be detected by various ground-based and space-based instruments. Non-detection of these Hawking emitted particles has been used to place significant constraints on their abundance [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54] (see also recent reviews on this topic [55, 56, 57, 58]). Beyond direct detection, the energy deposited into the IGM by Hawking radiation also leaves indirect imprints that can be probed through cosmological observations, including the cosmic microwave background (CMB) [59, 60, 61, 62, 63] and the global 21-cm signal [64, 65, 66, 67, 68]. Recently, Ref. [40] derived constraints on the abundance of light PBHs using IGM temperature measurements from the Lyman-$\alpha$ forest. In this paper, we significantly improve and extend their argument (see Section V) — arriving at an order of magnitude stronger constraint, which is more reliable and includes more detailed microphysics.
@@ II Hawking Radiation from Evaporating Primordial Black Holes
Non-rotating and uncharged PBHs of mass $M_{\mathrm{PBH}}$ have a temperature that depends solely on their mass [69, 70, 71, 72]
@@
where $\hbar=h/2\pi$ is the reduced Planck’s constant, $G_{\mathrm{N}}$ denotes the gravitational constant. These PBHs can emit Standard Model (SM) particles, such as photons, electrons, positrons etc, via Hawking radiation, gradually shedding their mass. The spectrum of these emitted particles can be parameterized as [69, 70, 71, 72]
@@
where $k_{\mathrm{B}}$ is the Boltzmann constant, $E$ is the total energy of the emitted particles, $s$ is the spin of the emitted particle, and $\mu$ denotes its rest mass. $\Gamma(E,\mu,T_{\mathrm{PBH}})$ denotes the dimensionless gray-body factor, representing the probability of an emitted particle escaping the black hole’s potential, and is evaluated numerically. We use the publicly available code BlackHawk [73] to calculate the primary Hawking emission spectra for various particles, and in Figure 1, we illustrate the emitted electron, positron, and photon spectra for a PBH of mass $7\times 10^{16}\,\mathrm{g}$. We do not show the neutrino spectrum, as the effect of neutrinos on the IGM is negligible. We also neglect potential secondary contributions in order to be conservative.
@@
where $f_{\mathrm{PBH}}$ is the fraction of dark matter composed of PBHs, $\rho_{c}$ is the critical density, and $\Omega_{\mathrm{DM}}$ is the dark matter density parameter [74].
@@ III IGM Thermal and Ionization History in Presence of PBH Energy Injection
Injected high-energy particles cool rapidly via various processes [75]. Electrons and positrons cool through inverse Compton scattering with the CMB and atomic processes like collisional ionization and excitation of hydrogen and helium atoms, and Coulomb heating of free electrons and ions in the IGM, while photons cool through redshifting, pair production, photon-photon scattering, Compton scattering, and photoionization of the hydrogen and helium atoms in the IGM. Considering the timescales of these processes, the injected energy from the PBHs can be assumed to deposit into four channels: (a) hydrogen ionization, (b) helium ionization, (c) Lyman-$\alpha$ excitation, and (d) Coulomb heating. Following Ref. [76], we encapsulate this in a numerical factor $f_{\mathrm{Cnl}}(z,\textbf{x})$, allowing the energy deposition rate into channel ‘Cnl’ to be expressed as
@@
where $\dot{T}_{\mathrm{adia}}$ represents the cooling due to adiabatic expansion, and $\dot{T}_{\mathrm{C}}$ accounts for heating or cooling due to (inverse) Compton scattering off the CMB. The term $\dot{T}_{\mathrm{atom}}$ encompasses all atomic cooling processes such as collisional ionization, collisional excitation, recombination, and Bremsstrahlung cooling. (Details on these processes are provided in Appendix A.) The term $\dot{T}_{\mathrm{reion}}$ includes the effect of photoheating during reionization, driven by active galactic nuclei (AGNs) and star-forming galaxies. To compute $\dot{T}_{\mathrm{reion}}$, we adopt the reionization model of Puchwein et al. [77], which self-consistently accounts for the evolution of galaxies and AGN, the resulting ultraviolet background (UVB), and its impact on the IGM by solving the radiative transfer equation. This approach yields a reionization and thermal history that is consistent with observational constraints as well as three-dimensional radiative transfer simulations of reionization. Puchwein et al. also demonstrate that applying their photoheating and photoionization rates to a gas cell at mean density, as we do in this work, accurately reproduces the temperature and ionization evolution predicted by full cosmological simulations (see their Figure C1). Additionally, we apply a uniform rescaling factor of 0.9 to the H i, He i, and He ii photoheating rates in the Puchwein et al. model to better match the observed IGM temperature. Finally, $\dot{T}_{\mathrm{PBH}}$, denotes the contribution of PBH energy injection to the temperature evolution of the IGM, and is given by
@@
where $\dot{x}_{\text{HII/HeII}}^{\mathrm{atom}}$ accounts for collisional ionization and recombination (see Appendix A for details), and $\dot{x}_{\text{HII/HeII}}^{\text{reion}}$ corresponds to the photoionization rates from the reionization model [77]. The contribution from PBHs to hydrogen ionization, is given by
@@
with collisional ionization and recombination coefficients taken from Ref. [78]. The contribution from reionization is
@@
The photoionization rates used in this calculation are taken from Ref. [77].
To compute the IGM thermal history in the presence of PBHs, we now need to solve the equations for $f_{\mathrm{Cnl}}$, $T_{\mathrm{IGM}}$, $x_{\mathrm{HII}}$, $x_{\mathrm{HeII}}$, and $x_{\mathrm{HeIII}}$ self-consistently. We do this using the publicly available code DarkHistory [76], by modifying the code to solve these equations down to $z=0$, and incorporating He iii evolution as well as PBH energy injection. We now describe these modifications to DarkHistory.
[…]

### v2

@@ Abstract
Primordial black holes (PBHs) have been explored as potential dark matter candidates, with various astrophysical observations placing upper limits on the fraction $f_{\mathrm{PBH}}$ of dark matter in the form of PBHs. However, a largely underutilized probe of PBH abundance is the temperature of the intergalactic medium (IGM), inferred from the thermal broadening of absorption lines in the Lyman-$\alpha$ forest of quasar spectra. PBHs inject energy into the IGM via Hawking radiation, altering its thermal evolution. In this work, we constrain this energy injection by self-consistently modeling its interplay with the cosmological ultraviolet background from galaxies and supermassive black holes. Leveraging IGM temperature measurements spanning the past twelve billion years ($z\sim 0$ to $6$), we derive one of the most stringent constraints on PBH-induced heating from light PBHs within the mass range $10^{15}$$10^{17}$ g. Specifically, for $M_{\mathrm{PBH}}=10^{16}$ g, we find $f_{\mathrm{PBH}}<5\times 10^{-5}$ at 95% confidence, with the bound scaling approximately as $M_{\mathrm{PBH}}^{4}$ at other masses. Our inclusion of helium reionization and low-redshift temperature measurements strengthens previous IGM-based PBH constraints by an order of magnitude or more. Compared to other existing limits, our result is among the strongest, second only to the constraints from the 511 keV line from the Galactic Center, but with distinct systematics. More broadly, this study highlights the IGM thermal history as a powerful and independent probe of beyond-standard-model physics.
@@ I Introduction
It is now known that the IGM at redshifts $z<6$ is in thermal equilibrium with a cosmological radiation background created by star-forming galaxies and accreting supermassive black holes. The balance between photoheating from this radiative background, heating from gravitational collapse, and cooling from Hubble expansion imparts to the IGM a power-law temperature–density relation, usually written as [1997MNRAS.292...27H],
@@
Fortunately, the temperature of the IGM can also be measured out at least to $z\sim 6$. Most of these measurements use observations of the Lyman-$\alpha$ forest to infer the IGM temperature. The Lyman-$\alpha$ forest is a series of absorption lines observed in the spectra of distant luminous quasars and serves as an excellent tracer of the intervening diffuse gas (see Refs. [2009RvMP...81.1405M, 2016ARA&A..54..313M] for reviews). The individual absorption lines in the Lyman-$\alpha$ forest are caused by the neutral hydrogen residing in the IGM. Consequently, the widths of these lines are sensitive to the thermal state of the IGM. The thermal motion of neutral hydrogen introduces a broadening in the width of the lines, expressed as $\Delta\nu=\nu_{\mathrm{Ly}\alpha}b_{\mathrm{th}}/c$, where $c$ is the speed of light, $\nu_{\mathrm{Ly}\alpha}$ is the Lyman-$\alpha$ absorption frequency, and $b_{\mathrm{th}}=\sqrt{2k_{\mathrm{B}}T/m_{\mathrm{H}}}$ is the Doppler parameter associated with the thermal motion with $k_{\mathrm{B}}$ being the Boltzmann’s constant, $m_{\mathrm{H}}$ the mass of the hydrogen atom, and $T$ the temperature of the IGM. By directly measuring the Doppler parameter of individual absorption lines in the Lyman-$\alpha$ forest, or by indirectly measuring it from other statistics sensitive to thermal broadening, it is therefore possible to infer the temperature of the IGM [2000MNRAS.318..817S, 2000ApJ...534...41R, 2001ApJ...557..519Z, 2002ApJ...567L.103T, 2010ApJ...718..199L, 2011MNRAS.410.1096B, 2015ApJ...799..196L, 2017PhLB..773..258G, 2017Sci...356..418R, 2017MNRAS.466.2690R, 2018MNRAS.474.2871R, 2019ApJ...872..101B, 2019ApJ...872...13W, 2019ApJ...876...71H, 2019ApJ...876...31G, 2021MNRAS.506.4389G, 2020MNRAS.494.5091G, 2022ApJ...933...59V, 2025MNRAS.536....1H]. This approach has revealed that $T_{\mathrm{IGM}}$ is of the order of $\sim 10^{4}$ K in the twelve billion years since $z\sim 6$, with some evolution. At $z\sim 6$, the temperature is around $11,500$ K at $z\sim 5.8$ and gradually declines to $\sim 7,300$ K by $z\sim 4.6$. Due to helium reionization at $z\sim 3$, the temperature peaks at approximately $14,750$ K before cooling again to $\sim 4,700$ K at $z\sim 0.06$. The typical uncertainties in these measurements range from $10\%$ to $20\%$.
@@
Exotic energy injection processes can significantly affect the thermal and ionization history of the IGM. Examples include dark matter decay or annihilation [Slatyer:2015jla, Slatyer:2015kla, Liu:2016cnk], dark matter-baryon interactions [Ali-Haimoud:2015pwa, Munoz:2015bca], ultralight dark photon dark matter [Mirizzi:2009iz, McDermott:2019lch], and light primordial black holes (PBHs) [Clark:2016nst, Mittal:2021egv]. Nonetheless, measurements of the IGM temperature from the Lyman-$\alpha$ forest as a probe of these processes have largely been overlooked so far. Indeed, while the Lyman-$\alpha$ forest has been used to constrain exotic processes, these constraints have primarily relied on measurements of the Lyman-$\alpha$ transmission power spectrum, which probes small-scale cosmological matter density fluctuations; these constraints have largely disregarded the information contained in the Lyman-$\alpha$ forest about the thermal state of the IGM. Examples include warm dark matter [2005PhRvD..71f3534V, 2008PhRvL.100d1304V, 2013PhRvD..88d3502V, 2017PhRvD..96b3522I, 2017JCAP...12..013B, 2020JCAP...04..038P, 2023PhRvD.108b3502V, 2023arXiv230904533I], sterile neutrinos [2005PhRvD..71f3534V, 2008PhRvL.100d1304V, 2017JCAP...12..013B, 2019MNRAS.489.3456G], fuzzy dark matter [2017PhRvL.119c1302I, 2017MNRAS.471.4606A], ultralight axions [2021PhRvL.126g1302R], primordial black holes (PBHs) with masses around 100100 M [2019PhRvL.123g1102M], and isocurvature fluctuations [2020PhRvD.101l3518I]. The potential of IGM temperature as a diagnostic tool for exotic physics remains largely untapped [2017JCAP...11..043M, 2021PhRvD.104d3514L, 2022PhRvL.129u1102B, 2024arXiv240910617S].
@@
In this study, we revisit the use of IGM temperature measurements from the Lyman-$\alpha$ forest to place constraints on light PBHs. Such light PBHs in the mass range of $10^{15}$ to $10^{17}\,\mathrm{g}$ have recently gained attention as potential dark matter candidates. Through Hawking evaporation, they emit particles that can be detected by various ground-based and space-based instruments. Non-detection of these Hawking emitted particles has been used to place significant constraints on their abundance [Carr:2009jm, Boudaud:2018hqb, Arbey:2019vqx, DeRocco:2019fjq, Laha:2019ssq, Dasgupta:2019cae, Laha:2020ivk, Chan:2020vsr, Ray:2021mxu, Coogan:2020tuf, Kim:2020ngi, Laha:2020vhg, Chan:2020zry, Ray:2022rfz, Bernal:2022swt, 2024JCAP...09..022T, 2024PhRvD.110l3022L] (see also recent reviews on this topic [Carr:2020gox, Carr:2020xqk, Green:2020jor, Escriva:2022duf]). Beyond direct detection, the energy deposited into the IGM by Hawking radiation also leaves indirect imprints that can be probed through cosmological observations, including the cosmic microwave background (CMB) [Clark:2016nst, Stocker:2018avm, Poulter:2019ooo, Acharya:2020jbv, Cang:2020aoo] and the global 21-cm signal [Clark:2018ghm, Mittal:2021egv, Natwariya:2021xki, Cang:2021owu, Saha:2021pqf]. Recently, Ref. [2024arXiv240910617S] derived constraints on the abundance of light PBHs using IGM temperature measurements from the Lyman-$\alpha$ forest. In this paper, we significantly improve and extend their argument (see Section V) — arriving at an order of magnitude stronger constraint, which is more reliable and includes more detailed microphysics.
@@ II Hawking Radiation from Evaporating Primordial Black Holes
Non-rotating and uncharged PBHs of mass $M_{\mathrm{PBH}}$ have a temperature that depends solely on their mass [Page:1976df, Page:1976ki, MacGibbon:1990zk, MacGibbon:2007yq]
@@
where $\hbar=h/2\pi$ is the reduced Planck’s constant, $G_{\mathrm{N}}$ denotes the gravitational constant. These PBHs can emit Standard Model (SM) particles, such as photons, electrons, positrons etc, via Hawking radiation, gradually shedding their mass. The spectrum of these emitted particles can be parametrized as [Page:1976df, Page:1976ki, MacGibbon:1990zk, MacGibbon:2007yq]
@@
where $k_{\mathrm{B}}$ is the Boltzmann constant, $E$ is the total energy of the emitted particles, $s$ is the spin of the emitted particle, and $\mu$ denotes its rest mass. $\Gamma(E,\mu,T_{\mathrm{PBH}})$ denotes the dimensionless gray-body factor, representing the probability of an emitted particle escaping the black hole’s potential, and is evaluated numerically. We use the publicly available code BlackHawk [2019EPJC...79..693A] to calculate the primary Hawking emission spectra for various particles, and in Figure 1, we illustrate the emitted electron, positron, and photon spectra for a PBH of mass $7\times 10^{16}\,\mathrm{g}$. We do not show the neutrino spectrum, as the effect of neutrinos on the IGM is negligible. We also neglect potential secondary contributions in order to be conservative.
@@
where $f_{\mathrm{PBH}}$ is the fraction of dark matter composed of PBHs, $\rho_{c}$ is the critical density, and $\Omega_{\mathrm{DM}}$ is the dark matter density parameter [Aghanim:2018eyx].
@@ III IGM Thermal and Ionization History in Presence of PBH Energy Injection
Injected high-energy particles cool rapidly via various processes [2009PhRvD..80d3526S]. Electrons and positrons cool through inverse Compton scattering with the CMB and atomic processes like collisional ionization and excitation of hydrogen and helium atoms, and Coulomb heating of free electrons and ions in the IGM, while photons cool through redshifting, pair production, photon-photon scattering, Compton scattering, and photoionization of the hydrogen and helium atoms in the IGM. Considering the timescales of these processes, the injected energy from the PBHs can be assumed to deposit into four channels: (a) hydrogen ionization, (b) helium ionization, (c) Lyman-$\alpha$ excitation, and (d) Coulomb heating. Following Ref. [2020PhRvD.101b3530L], we encapsulate this in a numerical factor $f_{\mathrm{Cnl}}(z,\textbf{x})$, allowing the energy deposition rate into channel ‘Cnl’ to be expressed as
@@
where $\dot{T}_{\mathrm{adia}}$ represents the cooling due to adiabatic expansion, and $\dot{T}_{\mathrm{C}}$ accounts for heating or cooling due to (inverse) Compton scattering off the CMB. The term $\dot{T}_{\mathrm{atom}}$ encompasses all atomic cooling processes such as collisional ionization, collisional excitation, recombination, and Bremsstrahlung cooling. (Details on these processes are provided in Appendix A.) The term $\dot{T}_{\mathrm{reion}}$ includes the effect of photoheating during reionization, driven by active galactic nuclei (AGNs) and star-forming galaxies. To compute $\dot{T}_{\mathrm{reion}}$, we adopt the reionization model of Puchwein et al. [2019MNRAS.485...47P], which self-consistently accounts for the evolution of galaxies and AGN, the resulting ultraviolet background (UVB), and its impact on the IGM by solving the radiative transfer equation. This approach yields a reionization and thermal history that is consistent with observational constraints as well as three-dimensional radiative transfer simulations of reionization. Puchwein et al. also demonstrate that applying their photoheating and photoionization rates to a gas cell at mean density, as we do in this work, accurately reproduces the temperature and ionization evolution predicted by full cosmological simulations (see their Figure C1). Additionally, we apply a uniform rescaling factor of 0.9 to the H i, He i, and He ii photoheating rates in the Puchwein et al. model to better match the observed IGM temperature. As the spectral shape of the ionizing background remains uncertain in UVB models, such rescaling is physically motivated: changing the photoheating rates is effectively equivalent to modifying the hardness of the ionizing spectrum (e.g., see Table 4 of [2021MNRAS.506.4389G]). Finally, $\dot{T}_{\mathrm{PBH}}$, denotes the contribution of PBH energy injection to the temperature evolution of the IGM, and is given by
@@
where $\dot{x}_{\text{HII/HeII}}^{\mathrm{atom}}$ accounts for collisional ionization and recombination (see Appendix A for details), and $\dot{x}_{\text{HII/HeII}}^{\text{reion}}$ corresponds to the photoionization rates from the reionization model [2019MNRAS.485...47P]. The contribution from PBHs to hydrogen ionization, is given by
@@
with collisional ionization and recombination coefficients taken from Ref. [2007MNRAS.374..493B]. The contribution from reionization is
@@
The photoionization rates used in this calculation are taken from Ref. [2019MNRAS.485...47P].
To compute the IGM thermal history in the presence of PBHs, we now need to solve the equations for $f_{\mathrm{Cnl}}$, $T_{\mathrm{IGM}}$, $x_{\mathrm{HII}}$, $x_{\mathrm{HeII}}$, and $x_{\mathrm{HeIII}}$ self-consistently. We do this using the publicly available code DarkHistory [2020PhRvD.101b3530L], by modifying the code to solve these equations down to $z=0$, and incorporating He iii evolution as well as PBH energy injection. We now describe these modifications to DarkHistory.
Figure 3: Left panel shows the evolution of temperature of the IGM at mean density with redshift for $M_{\mathrm{PBH}}=7\times 10^{16}$ g, with three $f_{\mathrm{PBH}}$ values, 0 (no PBHs; blue curve), $5\times 10^{-3}$ (orange curve), and $5\times 10^{-2}$ (green curve), in comparison with various measurements from $z\sim 0$66. The right panel shows the evolution of the ionization fractions of hydrogen (solid curve), singly-ionized helium (dashed curve), and doubly-ionized helium (dot-dashed curve). These ionization fractions remain largely unaffected by the inclusion of PBHs; it is the IGM temperature that shows the more significant effect. The gray curves in both panels represent a model without He iii evolution for $f_{\mathrm{PBH}}=5\times 10^{-3}$, which leads to an unphysical temperature increase.